OpenStudy (anonymous):

a vehicle travels at a certain average speed for a distance of 63km and then travels a distance of 72km at an average speed of 6km/h more than its original speed. if it takes 3 hours to complete the total journey, what si its original average speed?

OpenStudy (anonymous):

63/v + 72/(v+6) = 3

Parth (parthkohli):

The vehicle travels those 63 kilometers, let's say, at \(x\) kilometers per hour. Then the next 72 kilometers would be at a speed of \(x + 6\). How would you express the time taken to travel those \(63\) kilometers in terms of \(x\)?

Parth (parthkohli):

And how would you express the time taken to travel the next \(72\) kilometers?

OpenStudy (anonymous):

63/x - 72/x+ 64 = 3 ...... wt after ths?

OpenStudy (anonymous):

???

OpenStudy (ranga):

63/v + 72/(v+6) = 3 multiply each term by v(v+6) 63(v+6) + 72(v) = 3v(v+6) If you simplify you will get a quadratic equation in v. And then you can solve for v.

OpenStudy (anonymous):

i got, ... -3x^2 + 52x + 378 = 0 , ... now how to slove ths by completing the square method ?

OpenStudy (anonymous):

any1 thr???